# auto correlation matrix ADC analog to digital converter

### v1

@@
Resolution Improvement in OFDM-based
Joint Communication and Sensing through Combined Tracking and Interpolation thanks: This work has received funding in part from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001899) and in part from the German Federal Ministry of Education and Research (BMBF) within the projects Open6GHub (grant agreement 16KISK010) and KOMSENS-6G (grant agreement 16KISK123).
@@ Abstract
We investigate a monostatic orthogonal frequency-division multiplexing (OFDM)-based joint communication and sensing (JCAS) system with multiple antennas for object tracking. The native resolution of OFDM sensing, and radar sensing in general, is limited by the observation time and bandwidth. In this work, we improve the resolution through interpolation methods and tracking algorithms. We verify the resolution enhancement by comparing the root mean squared error (RMSE) of the estimated range, velocity and angle and by comparing the mean Euclidean distance between the estimated and true position. We demonstrate how both a Kalman filter for tracking, and interpolation methods using zero-padding and the chirp Z-transform (CZT) improve the estimation error. We discuss the computational complexity of the different methods. We propose the KalmanCZT approach that combines tracking via Kalman filtering and interpolation via the CZT, resulting in a solution with flexible resolution that significantly improves the range RMSE.
@@
Joint Communication and Sensing, OFDM, Tracking, Kalman Filter, Interpolation, 6G
@@ I Introduction
JCAS is a technology to be integrated into the new 6G standard [1] that adds a sensor functionality to wireless communication devices. Different features of the environment can be sensed as the communication signals are reflected by passive objects resulting in multiple reflections of the same signal being received. These reflections can be identified and attributed to different objects in the environment, transforming devices for wireless communication into potential sensors. Different applications have been proposed including weather monitoring, network optimization, traffic assistance or drone detection. Additionally, an improvement in energy and spectral efficiency is expected compared to deploying separate radar and communication systems.
@@
OFDM is especially suitable as a waveform for 6G JCAS as it is a well established waveform for communication and has been extensively studied for application in radar [2]. However, precise radar estimates rely on sufficient observation time and a large signal bandwidth. Bandwidth remains a finite resource as different systems operate simultaneously and interference has to be prevented. In 6G, we expect a limited bandwidth that is primarily chosen for communication, while a sensing functionality would benefit from a larger bandwidth. This limits the sensing resolution. In particular, the limited range resolution contributes severely to the estimation error when localizing passive objects.
If we want to capture a trajectory in time, limited resolution makes a single estimate inaccurate. Methods such as tracking and interpolation offer more accurate estimates of the spatial environment. A well-designed interpolation approach improves the ability of the system to distinguish between closely spaced objects and increases the estimation precision, effectively increasing the detail of the sensed environment. Interpolation approaches for sensing include zero-padding (ZP), the CZT [3], or superresolution methods [4]. Tracking additionally helps to maintain consistent observations of moving targets, allowing for the prediction of their future positions, which is vital for navigation, collision avoidance, and beamforming. If positions follow a trajectory, we can use prediction to refine the estimates, as the positions along a trajectory are typically correlated. Numerous studies applying Kalman filters to object tracking are part of the radar literature, e.g. in FMCW radar [5]. Especially, angle of arrival tracking has been of interest, since it enables precise beam tracking [6]. True position tracking can potentially further enhance energy-aware beam tracking. Additionally, object tracking is a basic task for OFDM-based JCAS systems and therefore of interest [7] in order to track, e.g., obstacles in traffic or drones. In order to enable real-time tracking, the complexity of the resolution enhancement algorithm needs to be sufficiently low, yet it should yield a position estimate that is as accurate as possible.
This paper specifically compares Kalman filtering structures and interpolation methods to improve tracking resolution in an OFDM-based JCAS system. We discuss how the computational complexity can be limited and how the estimation results are affected by interpolation and tracking, respectively. We present an alternative event-based measurement approach and propose the KalmanCZT as a combination approach of interpolation and tracking.
@@ II System Model
In this paper, we investigate a monostatic OFDM-based JCAS setup, where the transmitter and the sensing receiver are co-located, i.e., part of the same base station, and have multiple antennas. Our goal is to track the angle of arrival (AoA), range and radial velocity of an object based on its reflection to track its location in a 2D plane. The transmit signal is simultaneously used to communicate with a user equipment (UE) equipped with a single antenna in a separate area of interest. We simulate a split transmit beam to enable communication, yet focus on the object tracking in this work.
@@ II-A Transmitter
where $T_{0}=T_{\text{cp}}+T$ and $g(t)=1$ for $t\in[-T_{\text{cp}},T]$ and 00 elsewhere is a rectangular pulse shaping filter. We transmit the OFDM signal at carrier frequency $f_{\text{c}}$. The transmitter is equipped with $K$ antennas in a uniform linear array (ULA) with antenna spacing $c_{0}/2f_{\text{c}}$, where $c_{0}$ represents the speed of light. Beam squint is neglected in this configuration. After beamforming with the beamforming function $w_{a}(\varphi)$, with $\varphi$ being the AoA of the tracked object and antenna index $a\in[0,K\nobreak-1]$, the signal transmitted towards angle $\varphi$ can be expressed as
$\boldsymbol{s}_{\varphi}(t)=\sum_{a=0}^{K-1}w_{a}(\varphi)s(t)\mathrm{e}^{\mathrm{j}\pi a\sin\varphi}.$ (1)
@@ II-B Channels
A part of the transmit signal reaches the communication receiver while another part is reflected by the object of interest and reaches the sensing receiver co-located with the transmitter. The sensing channel is assumed to have one reflection caused by the tracked object with delay $\tau_{\mathrm{o}}$, [^1] We do not distinguish between random variables and scalars notation-wise to retain a clear and simple notation associated Doppler shift $f_{\mathrm{D},{\mathrm{o}}}$ and factor $a_{\mathrm{o}}=1$ representing the radar cross section, path loss and directivity gains. The received signal is given by
@@ II-C Sensing Receiver
At the sensing receiver, we first estimate the angle using the Bartlett method due to its computational complexity advantage in favor of more sophisticated methods [8]. Then, we perform receive beamforming based on the estimated angle $\varphi_{\text{est}}$. The received signal that is considered for estimating the radial velocity and the range is
@@
We can define the instantaneous signal-to-noise ratio (SNR) for AoA estimation to $\text{SNR}_{\varphi}=\sigma_{\text{s}}^{2}/\sigma_{\text{ns}}^{2}$, with $\sigma_{\text{s}}^{2}$ representing the end-to-end received signal power after effects such as path loss or the radar cross section.
@@
with $M\rightarrow$ denoting that the given operation is performed on each row of the input matrix and $N\downarrow$ denoting that the operation is performed on each column of the matrix. We estimate the range and velocity as the values $\hat{r}$ and $\hat{v}$ corresponding to the element of the RDM with maximum power, since we assume that there is always one object present. In Fig. 1, an example trajectory and the resulting estimated positions are shown.
@@ III Sensing Signal Processing
The computational complexity of the RDM can be inferred from the fast Fourier transform (FFT) complexity as $5NM\cdot\nobreak\log_{2}(NM)$ floating point operations [10, Table 2]. The signal resolution can be improved through interpolation methods, including ZP of the frame or applying CZTs, or by using time correlation in the positions and applying tracking algorithms. In this section, we will introduce these methods in detail.
@@ III-A Interpolation Methods
The application of ZP before applying the inverse fast Fourier transform (IFFT) and the FFT in the sensing processing (4) is equivalent to a sinc-interpolation of the RDM. Both the required memory for the transform result as well as the computational effort for calculating the interpolated RDM increase with the ZP amount. As we are mostly interested in improving the range resolution, we zero-pad the frame by adding rows to the input matrix resulting in a frame of dimensions $N_{\text{pad}}\times M$. After ZP, we need to calculate $N_{\text{pad}}$ IFFT values from $N$ nonzero entries per original column, while the complexity of the FFT per row doesn’t change. Therefore, the total complexity can be approximated as $5MN_{\text{pad}}\log_{2}(N_{\text{pad}})+5NM\log_{2}(M)$ FLOPs.
@@
The CZT generalizes the FFT to frequency components beyond the unit circle [11] and has been used to improve resolution in radar [12]. In our application, we are interested in the zoom capabilities of the transform, i.e., we use it to analyze a smaller section of the RDM in more detail. We first need a good estimate of the area where the tracked object is located. We can get this information by either using the last estimate or a prediction based on past dynamics as the center of an observation window, or by detecting with a lower resolution before refining our estimate. The CZT as well as its inverse can be calculated with a complexity of $75(2N\log_{2}(2N))$ FLOPs, assuming the number of output and input values is both $N$. Note that this is five times higher than the complexity reported for the CZT in [10, Table 2]. In [10], the CZT complexity is based only on the complexity of the underlying FFTs. However, we observe that the costs of other operations such as exponentiation are also significant. The total resulting complexity is $75M(2N\log_{2}(2N))+5NM\log_{2}(M)$ FLOPs. In our application, we leverage the RDM estimate in the first time step and then for the subsequent steps, we zoom around the estimate of the previous time step. The resulting transform to generate a zoomed in section is calculated by:
$A$ $=i_{\text{est}}/N_{\text{FFT}}-d/2,\quad W=\mathrm{e}^{\left(\mathrm{j}\frac{2\pi d}{N_{\text{CZT}}}\right)}$ (7)
@@
while $d$ invokes the index span that the zoomed transform is considering and $i_{\text{est}}$ being the index in the middle of the zoomed frame. From the elements of $\mathrm{RDM}_{\text{fine}}$ with maximum power, the corresponding range and velocity estimates can be obtained. These are more precise than estimation using the native RDM.
In our scenario, we use the CZT to increase the range resolution and use the RDM to obtain the velocity estimate.
@@ III-B1 Kalman Filter
The Kalman filter can be divided into two distinct phases: the prediction phase and the update phase. In the prediction phase, the filter uses the current estimate $\boldsymbol{x}_{k-1}$ and the system dynamics ${\boldsymbol{F}}$ to predict the range, velocity and angle at the next time step. With our model, the prediction equation of the Kalman filter in each time step $k$ is given by
$\boldsymbol{x}_{k|k-1}=\begin{pmatrix}r_{k|k-1}\\ v_{k|k-1}\\ \varphi_{k|k-1}\end{pmatrix}=\underbrace{\begin{pmatrix}1&-T_{0}&0\\ 0&1&0\\ 0&0&1\end{pmatrix}}_{{\boldsymbol{F}}}\cdot\boldsymbol{x}_{k-1},$ (9)
with $\boldsymbol{x}_{k-1}$ denoting the output estimate of the Kalman filter in time step $k-1$ and $r_{k|k-1}$, $v_{k|k-1}$, $\varphi_{k|k-1}$ denoting the predicted range, radial velocity and AoA, respectively. As the radial velocity is defined to be positive for movements towards the receiver, we predict the new range as $r_{k|k-1}=r_{k-1}-T_{0}v_{k-1}$. During the update phase, the filter incorporates new measurements to correct the predicted state, weighting the impact of the measurement with Kalman gain ${\boldsymbol{K}}$. By combining these two phases, the Kalman filter continuously refines its estimates, leading to an increased accuracy. The behavior of the system is influenced by the observation noise ${\boldsymbol{\Sigma}}_{\text{meas},k}$, modeling the noise caused by the resolution limit and the prediction noise ${\boldsymbol{\Sigma}}_{\text{pred},k}$. The computational complexity of a Kalman filter is $\mathcal{O}_{\text{KF}}(s^{3})$ with $s$ denoting the size of the state space (which is $3$ in our case).
III-C Event-Based Measurement (EBM)
Seeking a low-complexity solution for accurate object tracking and recognizing that the RDM estimates $\hat{r}$ and $\hat{v}$ provide limited information, we propose an alternative range estimation method that combines linear prediction with event-based range measurements:
$r_{\text{ebm},k}=\begin{cases}\frac{1}{2}(\hat{r}_{k}+\hat{r}_{k-1})&,\quad\hat{r}_{k}\neq\hat{r}_{k-1}\\ r_{\text{ebm},k-1}-T_{0}\hat{v}_{k}&,\quad\hat{r}_{k}=\hat{r}_{k-1}\end{cases}$ (10)
@@ III-C Event-Based Measurement (EBM)
Figure 2: Flowgraph of the novel, proposed joint tracking and filtering algorithm KalmanCZT
III-D KalmanCZT
We propose the KalmanCZT approach that combines tracking and filtering with a Kalman filter and the interpolation capabilities of the CZT. We specifically address improving the range resolution of object tracking. A high-level block diagram of the KalmanCZT is shown in Fig. 2. As the Kalman filter estimates the covariance matrix ${\boldsymbol{\Sigma}}_{\text{est},k|k-1}$ of the prediction, we use this estimate to steer the resolution of the CZT measurement, as well as using the predicted value $\boldsymbol{x}_{k|k-1}$ as the center of the observation window of the CZT.

### v4

@@
Improved Estimation Accuracy in OFDM-based
Joint Communication and Sensing through Kalman Tracking and Interpolation thanks: This work has received funding in part from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001899) and in part from the German Federal Ministry of Education and Research (BMBF) within the projects Open6GHub (grant agreement 16KISK010) and KOMSENS-6G (grant agreement 16KISK123).
@@ Abstract
We investigate a monostatic orthogonal frequency-division multiplexing (OFDM)-based joint communication and sensing (JCAS) system for object tracking. Our setup consists of a transmitter and receiver equipped with an antenna array for fully digital beamforming. The native resolution of range and velocity in all radar-like sensing, including OFDM radar sensing, is limited by the observation time and bandwidth. In this work, we improve the parameter estimates (estimates of range) through interpolation methods and tracking algorithms. We verify our method by comparing the root mean squared error (RMSE) of the estimated range, velocity and angle and by comparing the mean Euclidean distance between the estimated and true position. We demonstrate how both a Kalman filter for tracking, and interpolation methods using zero-padding and the chirp Z-transform (CZT) improve the estimation error. We discuss the computational complexity of the different methods. We propose the KalmanCZT approach that combines tracking via Kalman filtering and interpolation via the CZT, resulting in a solution with flexible resolution that significantly improves the range RMSE.
@@
Joint Communication and Sensing, OFDM, Tracking, Kalman Filter, Interpolation
@@ I Introduction
JCAS is a technology to be integrated into the new 6G standard [1] that adds a sensor functionality to wireless communication devices. Different features of the environment can be sensed as transmit signals are reflected by passive objects resulting in multiple reflections of the same signal being received. These reflections can be identified and attributed to different objects in the environment, transforming devices for wireless communication into potential sensors. Different applications have been proposed including weather monitoring, network optimization, traffic assistance or drone detection. Additionally, an improvement in energy and spectral efficiency is expected compared to deploying separate radar and communication systems.
@@
OFDM is especially suitable as a waveform for 6G JCAS as it is a well established waveform for communication, as it enables robust communication in frequency-selective channels for low computational resources. Additionally, it has been studied for application in radar for over a decade [2]. However, precise radar estimates rely on sufficient observation time and a large signal bandwidth. Bandwidth remains a finite resource as different systems operate simultaneously and interference has to be prevented. In 6G, we expect a limited bandwidth that is primarily chosen for communication, limiting the sensing resolution. In particular, the limited range resolution contributes severely to the estimation error when localizing objects that do not communicate.
If we want to capture a trajectory in time, limited resolution makes a single estimate inaccurate. Methods such as tracking and interpolation offer more accurate estimates of the spatial environment. A well-designed interpolation approach improves the ability of the system to distinguish between closely spaced objects and increases the estimation precision, effectively increasing the detail of the sensed environment. Interpolation approaches for sensing include zero-padding (ZP), the CZT [3], or super-resolution methods [4]. Tracking additionally helps to maintain consistent observations of moving targets, allowing for the prediction of their future positions, which is vital for navigation, collision avoidance, and beamforming. If positions follow a trajectory, we can use prediction to refine the estimates, as the positions along a trajectory are typically correlated. Numerous studies applying Kalman filters to object tracking are part of the radar literature, e.g. in FMCW radar [5]. Especially, angle of arrival tracking has been of interest, since it enables precise beam tracking [6]. True position tracking can potentially further enhance energy-aware beam tracking. Additionally, object tracking is a basic task for OFDM-based JCAS systems and therefore of interest [7] in order to track, e.g., obstacles in traffic or drones. In order to enable real-time tracking, the complexity of the resolution enhancement algorithm needs to be sufficiently low, yet it should yield a position estimate that is as accurate as possible.
This paper specifically compares Kalman filtering structures and interpolation methods to improve tracking estimation in an OFDM-based JCAS system. We discuss how the computational complexity can be limited and how the estimation results are affected by interpolation and tracking, respectively. We present an alternative event-based measurement approach and propose the KalmanCZT as a combination approach of interpolation and tracking.
@@ II System Model
In this paper, we investigate a monostatic OFDM-based JCAS setup, where the transmitter and the sensing receiver are co-located, i.e., part of the same base station, and are equipped with uniform linear arrays to perform digital beamforming. Our goal is to track the angle of arrival (AoA), range and radial velocity of a passive object based on its reflection to track its location in a 2D plane. The transmit signal is simultaneously used to communicate with a user equipment (UE) equipped with a single antenna in a separate area of interest. We simulate a split transmit beam to enable communication, yet focus on the object tracking in this work. This approach enables us to use samples beyond the pilots for sensing which increases the accuracy for sensing.
@@ II-A Transmitter
where $T_{0}=T_{\text{cp}}+T$ and $g(t)=1$ for $t\in[-T_{\text{cp}},T]$ and 00 elsewhere is a rectangular pulse shaping filter. We transmit the OFDM signal at carrier frequency $f_{\text{c}}$. The transmitter is equipped with $K$ antennas in a ULA with antenna spacing $c_{0}/2f_{\text{c}}$, where $c_{0}$ represents the speed of light. In this work, we assume a genie-aided beamformer to isolate the performance of the receiver. After beamforming with the beamforming function $\tilde{w}_{a}=\mathrm{e}^{-\mathrm{j}\pi a\sin\varphi}/2+\mathrm{e}^{-\mathrm{j}\pi a\sin\theta}/2$, with $\varphi$ being the angle of departure (AoD) of the tracked object, $\theta$ being the AoD of the communication receiver, and antenna index $a\in[0,K\nobreak-1]$, the signal transmitted towards angle $\varphi$ can be expressed as
$\boldsymbol{s}_{\varphi}(t)=\sum_{a=0}^{K-1}\tilde{w}_{a}s(t)\mathrm{e}^{\mathrm{j}\pi a\sin\varphi}.$ (1)
@@ II-B Channels
A part of the transmit signal is received by the UE while another part is reflected by the object of interest and reaches the sensing receiver co-located with the transmitter. The sensing channel is assumed to have one reflection caused by the tracked object with delay $\tau_{\mathrm{o}}$, [^1] We do not distinguish between random variables and scalars notation-wise to maintain readability associated Doppler shift $f_{\mathrm{D},{\mathrm{o}}}$ and factor $a_{\mathrm{o}}=1$ representing the radar cross section, path loss and directivity gains. The received signal is given by
@@ II-C Sensing Receiver
At the sensing receiver, we first estimate the angle $\varphi_{\text{est}}$ using the Bartlett method due to its computational complexity advantage [8]. Then, we perform receive beamforming based on $\varphi_{\text{est}}$. The received signal that is considered for estimating the radial velocity and the range, with $w_{a}=\mathrm{e}^{-\mathrm{j}\pi a\sin\varphi_{\text{est}}}$, is
@@
We can define the instantaneous signal-to-noise ratio (SNR) for AoA estimation to $\text{SNR}_{\varphi}=\sigma_{\text{s}}^{2}/\sigma_{\text{ns}}^{2}$, with $\sigma_{\text{s}}^{2}$ representing the end-to-end received signal power.
@@
with $M\rightarrow$ denoting that the given operation is performed on each row of the input matrix and $N\downarrow$ denoting that the operation is performed on each column of the matrix. We estimate the range and velocity as the values $\hat{r}$ and $\hat{v}$ corresponding to the element of the RDM with maximum power, since we assume that there is always one object present. This peak detection is used for all methods described in this work, a discussion of different detectors such as super-resolution methods is left for future work due to the difficulty in computing their exact computation complexity. In Fig. 1, an example trajectory and the resulting estimated positions are shown.
@@ III Sensing Signal Processing
The complexity of computing the RDM can be inferred from the fast Fourier transform (FFT) complexity as $5NM\cdot\nobreak\log_{2}(NM)$ floating point operations [10, Table 2]. Signal estimation can be improved through interpolation methods, including ZP of the frame or applying CZTs, or by using time correlation in the positions and applying tracking algorithms. In this section, we will introduce these methods in detail.
@@ III-A Interpolation Methods
The application of ZP before applying the inverse fast Fourier transform (IFFT) and the FFT in the sensing processing (4) is equivalent to a sinc-interpolation of the RDM. Both the required memory for the transform result as well as the computational effort for calculating the interpolated RDM increase with the ZP amount. As we are mostly interested in improving the range estimation, we zero-pad the frame by adding rows to the input matrix resulting in a frame of dimensions $N_{\text{pad}}\times M$. After ZP, we need to calculate $N_{\text{pad}}$ IFFT values from $N$ nonzero entries per original column, while the complexity of the FFT per row doesn’t change. Therefore, the total complexity can be approximated as $5MN_{\text{pad}}\log_{2}(N_{\text{pad}})+5NM\log_{2}(M)$ FLOPs.
@@
The CZT generalizes the FFT to frequency components beyond the unit circle [11] and has been used to improve estimation in radar [12]. In our application, we are interested in the zoom capabilities of the transform, i.e., we use it to analyze a smaller section of the RDM in more detail. In the initial estimate, we use the whole RDM as an observation window as we assume no knowledge of the initial position. In the subsequent steps, we use the last estimate as the center of the observation window. The CZT as well as its inverse can be calculated with a complexity of $75(2N\log_{2}(2N))$ FLOPs, assuming the number of output and input values is both $N$. Note that this is five times higher than the complexity reported for the CZT in [10, Table 2]. In [10], the CZT complexity is based only on the complexity of the underlying FFTs. However, we observe that the costs of other operations such as exponentiation are also significant. The total resulting complexity is $75M(2N\log_{2}(2N))+5NM\log_{2}(M)$ FLOPs. The resulting transform to generate a zoomed in section is calculated using:
@@
where the parameters are chosen as $W_{\text{r}}=\mathrm{e}^{\mathrm{j}2\pi r_{\text{CZT}}/(N^{2}r_{\text{res}})}$ $A_{\text{r}}=\mathrm{e}^{\mathrm{j}2\pi\left(\frac{\hat{r}}{Nr_{\text{res}}}-\frac{r_{\text{CZT}}}{2Nr_{\text{res}}}\right)}$, $W_{\text{v}}=\mathrm{e}^{\mathrm{j}2\pi/M}$ and $A_{\text{v}}=-1$, with the last range estimate $\hat{r}$ as the center of the range observation window of size $r_{\text{CZT}}$. The range and velocity estimates corresponding to the element of $\mathrm{RDM}_{\text{fine}}$ with maximum power can then be obtained. These are more precise than estimation using the native RDM.
In our scenario, we use the CZT to increase the range estimation accuracy.
@@ III-B1 Kalman Filter
The Kalman filter can be divided into two distinct phases: the prediction phase and the update phase. In the prediction phase, the filter uses the current estimate $\boldsymbol{x}_{k-1}$ and the system dynamics ${\boldsymbol{F}}$ to predict the range, velocity and angle at the next time step. With our model, the prediction equation of the Kalman filter at time step $k$ is given by
$\boldsymbol{x}_{k|k-1}=\begin{pmatrix}r_{k|k-1}\\ v_{k|k-1}\\ \varphi_{k|k-1}\end{pmatrix}=\underbrace{\begin{pmatrix}1&-MT_{0}&0\\ 0&1&0\\ 0&0&1\end{pmatrix}}_{{\boldsymbol{F}}}\cdot\boldsymbol{x}_{k-1},$ (8)
with $\boldsymbol{x}_{k-1}$ denoting the output estimate of the Kalman filter in time step $k-1$ and $r_{k|k-1}$, $v_{k|k-1}$, $\varphi_{k|k-1}$ denoting the predicted range, radial velocity and AoA, respectively. As the radial velocity is defined to be positive for movements towards the receiver, we predict the new range as $r_{k|k-1}=r_{k-1}-MT_{0}v_{k-1}$. During the update phase, the filter incorporates new measurements to correct the predicted state, weighting the impact of the measurement with Kalman gain ${\boldsymbol{K}}$. By combining these two phases, the Kalman filter continuously refines its estimates, leading to an increased accuracy. The behavior of the system is influenced by the observation noise ${\boldsymbol{\Sigma}}_{\text{meas},k}$, the noise caused by the resolution limit and the prediction noise ${\boldsymbol{\Sigma}}_{\text{pred},k}$. The computational complexity of a Kalman filter is $\mathcal{O}_{\text{KF}}(s^{3})$ with $s$ denoting the size of the state space (which is $3$ in our case).
III-B2 Event-Based Measurement (EBM)
Seeking a low-complexity solution for accurate object tracking and recognizing that the RDM estimates $\hat{r}$ and $\hat{v}$ provide limited information, we propose an alternative range estimation method that combines linear prediction with event-based measurements:
$r_{\text{ebm},k}=\begin{cases}\frac{1}{2}(\hat{r}_{k}+\hat{r}_{k-1})&,\quad\hat{r}_{k}\neq\hat{r}_{k-1}\\ r_{\text{ebm},k-1}-MT_{0}\hat{v}_{k}&,\quad\hat{r}_{k}=\hat{r}_{k-1}\end{cases}$ (9)
@@ III-B2 Event-Based Measurement (EBM)
Figure 2: Flowgraph of the proposed joint tracking and filtering algorithm KalmanCZT
III-C KalmanCZT
We propose the KalmanCZT approach that combines tracking and filtering with a Kalman filter and the interpolation capabilities of the CZT. We specifically address improving the range estimation of object tracking. A high-level block diagram of the KalmanCZT is shown in Fig. 2. As the Kalman filter estimates the covariance matrix ${\boldsymbol{\Sigma}}_{\text{est},k|k-1}$ of the prediction, we use this estimate to determine the observation window of the CZT, as well as using the predicted value $\boldsymbol{x}_{k|k-1}$ as the center of the observation window of the CZT.